Bridgeport Ltd. estimates sales for the second quarter of 2017 will be as follows.

Month

Units

April 2,520
May 2,460
June 2,360


The target ending inventory of finished products is as follows.

March 31 2,050
April 30 2,230
May 31 2,100
June 30 2,300


Two units of material are required for each unit of finished product. Production for July is estimated at 2,690 units to start building inventory for the fall sales period. Bridgeport’s policy is to have an inventory of raw materials at the end of each month equal to 50% of the following month’s production requirements.

Raw materials are expected to cost $6 per unit throughout the period.

Calculate the May raw materials purchases in dollars.

Raw material purchases cost $
Answer #1

Raw material required in May = (2,460 + 2,100 - 2,230) X 2 = 4,660

Out of 4,660, 50% is already maintained in opening raw material.

= 4,660 - 50%

= 2,330 (A)

Closing Inventory = (2,360 + 2,300 - 2,100) X 2 X 50% = 2,560 (B)

(A) + (B) = 2,330 + 2,560 = 4,890

Raw material purchases = 4,890 X $6 = $29,340